OpenStudy (anonymous):

A curve is such athat {dy/dx}= {12 over (2x+1)^2} and P(1,5) is a point on the curve. Um... the last part of the question says, A point is moving along the curve in such a way that the x=coordinate is increasing at a constant rate of 0.3 units per second. Find the rate of increase of the y-coordinate when x=1

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

the dy/dx value at x=1

OpenStudy (experimentx):

dy = f'(1)x0.3

OpenStudy (anonymous):

sorry, 4/3 * 0.3= 0.4

OpenStudy (anonymous):

So, how?

OpenStudy (anonymous):

dy/dx value is the ratio of rate of change of y wrt rate of change of x. dy/dx at x=1 is 12/3^2=4/3 rate of change of x= 0.3 so, rate of change of y=4/3 * 0.3= 0.4

OpenStudy (anonymous):

thanks!